OpenStudy (anonymous):

What is the sum of a 30-term arithmetic sequence where the first term is 71 and the last term is -103? Answers to choose from. a)-480 b)-448 c)-416 d)-384

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

use reqd sum=(n/2)(first term +last term)=(30/2)(71+ -103) = -480

OpenStudy (anonymous):

Thank you soo much i have 2 more questions left to do then i can go to bed would you please help me get them done ill give you medals for both? @matricked

OpenStudy (anonymous):

if possible i will help u...

OpenStudy (anonymous):

Thank you soo soo much

OpenStudy (anonymous):

welcome